OpenStudy (anonymous):

When the positive integer k is divided by 7, the reminder is 6. What is the remainder when k+2 is divided by 7? (A) 0 (B) 1 (C) 2 (D) 3 (E) 4

OpenStudy (anonymous):

Somebody Please Help!!

OpenStudy (anonymous):

take an example 13 divided by 7 = 1 remainder 6 add 2 to 13 gives 15 15 divied by 7 whats the remainder

OpenStudy (anonymous):

remainder is 1

OpenStudy (anonymous):

right - and this is true for any number giving a remainder of 6 when divided by 7

OpenStudy (anonymous):

thanks Bro

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you can write the number rem 6 like k = 7n + 6 where n is a poistive integer then k + 2 = 7n + 6 + 2 = 7n + 7 + 1 = 7(n + 1) + 1 so remainder is 1